OpenStudy (anonymous):

Write an appropriate inverse variation equation if y = 9 when x = 4.

OpenStudy (anonymous):

let the reqd equation be of the form y = k/x we just need to know k the constant of proportionality

OpenStudy (anonymous):

thus k=xy as y = 9 when x = 4. k=9*4 =36 hence eq becomes y=36/x

OpenStudy (anonymous):

I almost had it! Omg thank you so much, Tough problem answerd by a great person. I thought it was 46

OpenStudy (whpalmer4):

Here's a handy list of how to construct the variation equations: "x varies directly with y" => y = kx "x varies inversely with y" => y = k/x "x varies jointly with y and z" => x = kyz in all cases, there's a k, though the value of k may be 1.

OpenStudy (whpalmer4):

and if you see something like "y varies inversely with x squared" you just plug that in, too: y = k/x^2